OpenStudy (anonymous):

Find the linearization of f(x) =1/sqrt(x+1) at x=0

myininaya (myininaya):

you know how find tangent lines right this is what it is asking you to do find the tangent line of the curve at x=0

myininaya (myininaya):

if x=0, then y=1/sqrt{0+1}=1/1=1 so we know a point on the the tangent line which is (1,1) find f' and plug in 0 for x to find the slope of the tangent line so then you can plug everything you find into y-y_1=m(x-x_1)

OpenStudy (anonymous):

I got y=1/10sqrt(5)+1/5, but that not one of the answer choices

OpenStudy (anonymous):

I got y=1/10sqrt(5)x+1/5, but that not one of the answer choices

myininaya (myininaya):

\[f'(x)=\frac{-1}{2}(x+1)^\frac{-3}{2}=\frac{-1}{2(x+1)^\frac{3}{2}} =>f'(0)=\frac{-1}{2}\]

myininaya (myininaya):

\[y-1=\frac{-1}{2}(x-0)\]

OpenStudy (anonymous):

it should be x+5 in the parenthesis instead of x+1

myininaya (myininaya):

ok well i did it for y=1/sqrt{x+1} (<-what you wrote above?) but you can do it for y=1/sqrt{x+5} use what i did as an example

OpenStudy (anonymous):

I wrote the problem wrong it should be x+5 inside the square root, sorry!

OpenStudy (anonymous):

y-1/5=1/10sqrt(5)(x-0)

myininaya (myininaya):

what happen to the square root when you plugged in 0 for x?

OpenStudy (anonymous):

the 5 is still inside the square root

myininaya (myininaya):

\[f(0)=\frac{1}{\sqrt{0+5}}=\frac{1}{\sqrt{5}} \cdot \frac{\sqrt{5}}{\sqrt{5}}=\frac{\sqrt{5}}{5}\]

myininaya (myininaya):

\[y-f(0)=m(x-0)\]

myininaya (myininaya):

so we need m

myininaya (myininaya):

\[f'(x)=\frac{-1}{2}(x+5)^\frac{-3}{2}=\frac{-1}{2(x+5)^\frac{3}{2}} =>f'(0)=\frac{-1}{2(5)^\frac{3}{2}}\]

myininaya (myininaya):

m=f'(0)

OpenStudy (anonymous):

here are the answers choices

OpenStudy (anonymous):

it doesn't match the answer choices

myininaya (myininaya):

i see the answer

myininaya (myininaya):

try simplifying and writing it as y=mx+b by adding f(0) on both sides

OpenStudy (anonymous):

hello all @ta i wrote out an answer for your last problem

OpenStudy (anonymous):

in your previous post oh actually it was this problem! i wrote it out, it is probably still there

myininaya (myininaya):

write as y=b+mx then factor anything the b and m have in common and voolah

OpenStudy (anonymous):

of course i made my usual typo, but i fixed it